# Resonant Multi-Scalar Production in the Generic Complex Singlet Model in the Multi-TeV Region

### v1

@@ I Introduction
In the SM, the di-Higgs production from gluon fusion at the $\sqrt{S}=13$ TeV LHC is predicted to be $\sigma=31.05$ fb Dawson:2022zbb ; Dawson:1998py ; Borowka:2016ehy ; Baglio:2018lrj ; Grazzini:2018bsd ; Baglio:2020wgt . Due to its small rate, double Higgs production can be quite sensitive to new physics contributions DiMicco:2019ngk . Of particular interest are scalar extensions of the SM that can change the nature of EWSB and the Higgs self-couplings. Indeed, there has been interest investigating scenarios in which the first evidence of new physics could be through variations of the Higgs tri-linear couplings DiLuzio:2017tfn ; Chang:2019vez ; Bahl:2022jnx ; Durieux:2022hbu . Another spectacular signature of additional scalars could be resonant di-Higgs production. Such a signal is much studied and appears in even the simplest of all possible extensions: a real scalar that is a singlet under the SM gauge group Silveira:1985rk ; OConnell:2006rsp ; Barger:2007im ; Bowen:2007ia ; Profumo:2007wc ; No:2013wsa ; Pruna:2013bma ; Profumo:2014opa ; Chen:2014ask ; Dawson:2015haa ; Buttazzo:2015bka ; Robens:2015gla ; Robens:2016xkb ; Kotwal:2016tex ; Lewis:2017dme ; Huang:2017jws ; Dawson:2018dcd ; Li:2019tfd ; Alves:2020bpi ; Papaefstathiou:2020iag ; Dawson:2020oco ; Adhikari:2020vqo ; Papaefstathiou:2021glr ; Dawson:2021jcl .
@@
There has been recent interest is moving beyond di-Higgs production and study more generic di-scalar production. For example, non-resonant production of two new scalars can provide insight into the early universe electroweak phase transition Curtin:2014jma ; Chen:2017qcz . A more spectacular signal is asymmetric resonant production of the SM Higgs+a new scalar or symmetric resonant production of two new scalars Basler:2018dac ; Abouabid:2021yvw ; Adhikari:2022yaa ; Dawson:2017jja ; Costa:2015llh ; Robens:2019kga ; Papaefstathiou:2020lyp . In this paper, we develop benchmark points for such signals in the SM extended by a complex scalar that is a singlet under the SM gauge group Coimbra:2013qq ; Costa:2014qga ; Ferreira:2016tcu ; Muhlleitner:2017dkd ; Basler:2019nas ; Muhlleitner:2020wwk ; Egle:2022wmq ; Egle:2023pbm ; Barger:2008jx ; Alexander-Nunneley:2010tyr ; Dawson:2017jja ; Adhikari:2022yaa ; Costa:2015llh ; Robens:2019kga ; Papaefstathiou:2020lyp ; Dawson:2023oce . Beyond providing interesting collider signatures of new scalars, the complex singlet extension can also help provide a strong first order electroweak phase transition necessary to generate the baryon asymmetry of the universe via electroweak baryogenesis Barger:2008jx ; Chao:2014ina ; Jiang:2015cwa ; Chiang:2017nmu ; Cheng:2018ajh ; Grzadkowski:2018nbc ; Chen:2019ebq ; Chiang:2019oms ; Cho:2021itv ; Cho:2022our ; Cho:2022zfg ; Zhang:2023mnu . This extension has been much studied in the scenarios where there are additional $\mathbb{Z}_{2}$ or $U(1)$ symmetries McDonald:1993ex ; Cerdeno:2006ha ; Barger:2008jx ; Coimbra:2013qq ; Costa:2014qga ; Costa:2015llh ; Ferreira:2016tcu ; Muhlleitner:2017dkd ; Basler:2019nas ; Muhlleitner:2020wwk ; Robens:2019kga ; Papaefstathiou:2020lyp ; Egle:2022wmq ; Egle:2023pbm . We will focus on the scenario in which there is no additional symmetry beyond the SM gauge group Dawson:2017jja ; Adhikari:2022yaa .
@@
If the mixing between $h_{3}$ and the SM Higgs boson is minuscule, the coupling between $h_{3}$ and SM fermions and gauge bosons will be negligible. The major production mode of $h_{3}$ would then be directly through these decays. It is possible that searches for resonant production of $h_{2}\rightarrow h_{1}h_{3}$ and $h_{2}\rightarrow h_{3}h_{3}$ could provide a discovery of two new real scalars at once, and could be the only feasible mechanisms to discover $h_{3}$ Dawson:2017jja .
@@ II Model
$V(\Phi,S_{c})$ $=$ $-\frac{\mu^{2}}{2}\Phi^{\dagger}\Phi+\frac{\lambda}{4}(\Phi^{\dagger}\Phi)^{4}+\frac{b_{2}}{2}|S_{c}|^{2}+\frac{d_{2}}{4}|S_{c}|^{4}+\frac{\delta_{2}}{2}\Phi^{\dagger}\Phi|S_{c}|^{2}$ (2)
@@
where the mass eigenstates are $h_{1},\,h_{2},\,$ and $h_{3}$ with masses $m_{1},\,m_{2},\,$ and $m_{3}$, respectively. We will associate $h_{1}$ with the observed scalar and set $m_{1}=125$ GeV ATLAS:2022vkf ; CMS:2022dwd . There is in principle a third $SO(3)$ rotation angle in Eq. (4), but it can be removed via a phase rotation of $S_{c}\rightarrow e^{i\,\theta_{3}}S_{c}$ Dawson:2017jja . Such a phase rotation is at most a redefinition of parameters in the scalar potential [Eq. (2)]. Hence, the third rotation angle can be removed with no physical effect. The free parameters of the model are then
@@
By requiring that $(\langle\Phi^{\dagger}\Phi\rangle,\langle S_{c}\rangle)=(v_{EW}^{2}/2,0)$ be an extremum of the potential and going to the scalar mass basis, the Lagrangian parameters can be related to these free parameters according to the following equations:
@@
Note that the coefficient of Eq. (15) differs by a factor of 2 from the other partial widths due to the two final states scalars not being identical. The general expressions for the trilinears in Eq. (13) are complicated, but using Eqs. (2,12), the $\theta_{2}\rightarrow 0$ limit of the relevant trilinear couplings are
$\lambda_{112}$ $=$ $\sin\theta_{1}\frac{m_{2}^{2}}{v_{\rm EW}}\left[\cos^{2}\theta_{1}\left(1+2\frac{m_{1}^{2}}{m_{2}^{2}}\right)+\frac{1}{\sqrt{2}}\cos\theta_{1}\sin\theta_{1}\left(\textrm{Re}\left(e_{1}\right)+\textrm{Re}\left(e_{2}\right)\right)\frac{v_{EW}}{m_{2}^{2}}\right.$
$\left.-\left(1-\frac{3}{2}\sin^{2}\theta_{1}\right)\frac{v_{EW}^{2}}{m_{2}^{2}}\left(\delta_{2}+\textrm{Re}\left(\delta_{3}\right)\right)\right]$
@@ III.1.1 Boundedness and Global Minima
$\frac{\lambda}{4}(\Phi^{\dagger}\Phi)^{4}+\frac{d_{2}}{2}|S_{c}|^{4}+\frac{\delta_{2}}{2}\Phi^{\dagger}\Phi|S_{c}|^{2}+\left(\frac{\delta_{3}}{4}\Phi^{\dagger}\Phi S_{c}^{2}+\frac{d_{1}}{8}S^{4}_{c}+\frac{d_{3}}{8}S_{c}^{2}|S_{c}|^{2}+{\rm h.c.}\right)\geq 0.$ (22)
@@ III.1.2 Perturbative Unitarity
Notably, the bound on $\lambda$ in Eq. 26 places a constraint on the allowed masses and mixings. For a particular choice of mixing angles, this induces an upper limit on the allowed heavy Higgs masses. In practice, as with the boundedness and global minimum constraints, we check numerically that the eigenvalues of Eq. (25) are bounded by $8\pi$.
@@ III.2 Experimental Constraints
We now consider experimental constraints on this model. First, we take Higgs precision measurements into account ATLAS:2022vkf ; CMS:2022dwd ; ATLAS-CONF-2022-067 ; CMS-PAS-HIG-19-011 . These measurements are typically given in terms of signal strengths, i.e. ratios of BSM and SM predictions for different Higgs production and decay channels:
@@
Note that these conclusions are independent of the choice of the other mixing angle $\theta_{2}$. The mixing angle $\theta_{1}$ can then be fit using a $\chi^{2}$ distributions:
@@
where $\Gamma_{\rm SM}(h_{2})$ is the total SM-like width at the mass $m_{2}$. From this it can be shown that branching ratio of $h_{2}$ into SM final states is
@@
Traditionally, a parameter point is accepted if the predicted cross section is lower than all of the observed 95% upper limits on cross sections reported in searches for new particles. The point is rejected if the predicted cross section exceeds an observed cross section limit in any channel. For example, see Refs. Bechtle:2013xfa ; Bechtle:2020uwn . We refer to this as the “hard cut” method. The red dashed curves in Figs. 1(a,b,c) show the resulting limits on $\sin\theta_{1}$ from utilizing the hard cut method. The list of included experimental results are given in Table 1 in App. D. Since $h_{2}$ couples to SM fermions and gauge bosons like the SM Higgs except suppressed by $\sin\theta_{1}$, its production and decay are similar to a SM Higgs of the same mass, except with the additional decay to scalars is available. Therefore, its dominate decays will be to gauge bosons or scalars and we only consider those experimental limits.
Figure 1: The 95% CL upper limits on $|\sin\theta_{1}|$ for (a) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$, (b) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, (c) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ profiled over, and (d) a variety of ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$. In all plot (black solid) the bounds from Higgs precision are shown. In (a-c) the bounds resulting from the (red dashed) hard cuts on scalar searches, (magenta dot-dot-dash) $\Delta\chi^{2}$ scalar search fits from Eq. (40), and (blue solid) $\Delta\chi^{2}$ combination of scalar searches and Higgs precision from Eq. (41). The results in (d) are from the $\Delta\chi^{2}$ combination of scalar searches and Higgs precision.
@@
where $\sigma_{i}(pp\rightarrow h_{2})$ is the resonance production cross section from initial state $i$, ${\rm BR}(h_{2}\rightarrow f)$ is the branching ratio into final state $f$, $\hat{\sigma}_{i,Exp}^{f}$ ($\hat{\sigma}_{i,Obs}^{f}$) is the expected (observed) 95%CL upper limit on $\sigma(i\rightarrow h_{2}){\rm BR}(h_{2}\rightarrow f)$. It can be checked that for an individual search, this $\chi^{2}$ reproduces the limit $\sigma(i\rightarrow h_{2}){\rm BR}(h_{2}\rightarrow f)<\hat{\sigma}_{i,Obs}^{f}$, consistent with the hard cut method. With this definition, all heavy scalar search channels can be consistently combined:
@@
$\chi^{2}_{Tot}=\chi^{2}_{h_{1}}+\chi^{2}_{h_{2}}.$ (41)
The magenta dash-dot-dot lines in Figs. 1(a,b,c) show the 95% CL upper limits on $|\sin\theta_{1}|$ that results from fitting scalar searches according to Eq. (40). For all limits we have assumed ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})={\rm BR}(h_{2}\rightarrow h_{3}h_{3})=0$. In (a) we set ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$, in (b) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, and in (c) we profile over ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$. As can be clearly seen, the results of Eq. (40) have similar behavior to and are consistently stronger than the hard cuts. This can be understood by noting that since all production cross sections of $h_{2}$ are proportional to $\sin^{2}\theta_{1}$, any increase in $|\sin\theta_{1}|$ essentially increases all signal rates in all channels uniformly increasing the $\chi^{2}$ value. Accordingly, we would more-or-less expect to see a global upward fluctuation in the channels listed in Tab. 1 if the scalar resonance exists. Such a global increase should result in a stronger constraint. Our $\chi^{2}$ method accounts for this, while the hard cut method does not.
The result from combining the scalar searches and Higgs precision data from Eq. (41) are shown as blue solid lines in Fig. 1 (a,b,c). For ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, the combined $\chi^{2}_{Tot}$ fit gives very similar results as the scalar search $\chi^{2}_{h_{2}}$. For the results with ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$ or ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ profiled over, combining the Higgs precision constraints with the scalar searches strengthens the bounds in the large mass region: $m_{2}\gtrsim 800$ GeV. Indeed, although the scalar search constraints may get weaker than the Higgs precision at very high mass, the combined Higgs and scalar search results stay more or similarly stringent as the Higgs measurement results.
@@
There are also constraints on the singlet mixing angle from electroweak precision observables Lopez-Val:2014jva ; deBlas:2016ojx ; Dawson:2017jja ; Ilnicka:2018def . For small $\theta_{2}$, fits to the oblique parameters Dawson:2017jja ; Ilnicka:2018def are always weaker than the most stringent constraints in Fig. 1. The most strongest electroweak precision constraint comes from fitting the $W$-mass individually Lopez-Val:2014jva . This is most relevant in the high mass region where the $W$-mass limit is $|\sin\theta_{1}|\lesssim 0.20-0.21$ for $m_{2}\gtrsim 800$ GeV Lopez-Val:2014jva ; Dawson:2017jja . Again, the most stringent constraints in Fig. 1 are as strong or stronger than the $W$-mass limit. Hence, for the rest of the paper we consider constraints on $\sin\theta_{1}$ from direct searches and Higgs precision data only.
Figure 2: Maximum allowed resonant production rates for $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in (a) gluon fusion and (b) VBF channels for (red dashed) hard cut bounds, (magenta dot-dot-dash) $\Delta\chi^{2}$ bounds on scalar searches from Eq. (40), and (blue solid) $\Delta\chi^{2}$ bounds on scalar searches and Higgs precision from Eq. (41). For comparison, SM predictions for (black solid) $h_{2}$ production and (green dot-dash-dash) $h_{1}h_{1}$ production are shown.
In Fig. 2 we show 95% CL upper limits on the cross section for production and decay $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in both the (a) gluon fusion and (b) VBF channels. These bounds are found by performing a simultaneous fit to $\sin\theta_{1}$ and $\sin^{2}\theta_{1}{\rm BR}(h_{2}\rightarrow h_{1}h_{1}$) with ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})={\rm BR}(h_{2}\rightarrow h_{3}h_{3})=0$. For the $\Delta\chi^{2}$ fits, $\sin\theta_{1}$ is profiled over. For the hard cut, a value of $\sin^{2}\theta_{1}{\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ is accepted if there exists a $\sin\theta_{1}$ that passes all cuts and Higgs precision data. For comparison, the SM single $h_{2}$ and double $h_{1}h_{1}$ production rates are shown. The $\Delta\chi^{2}$ and hard cut bounds produce very similar limits on $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$. For $m_{2}\lesssim 500$ GeV, the resonant gluon fusion di-Higgs rate can still be above the SM di-Higgs rate rate, up to an order of magnitude large. For VBF, the allowed rate is above the SM VBF di-Higgs rate for all masses shown. The resonant VBF rate can be up to a factor of 20 larger than the SM VBF di-Higgs prediction.

### v2

@@ I Introduction
In the SM, the di-Higgs production from gluon fusion at the $\sqrt{S}=13$ TeV LHC is predicted to be $\sigma=31.05$ fb Dawson:2022zbb ; Dawson:1998py ; Borowka:2016ehy ; Baglio:2018lrj ; Grazzini:2018bsd ; Baglio:2020wgt . Due to its small rate, double Higgs production can be quite sensitive to new physics contributions DiMicco:2019ngk . Of particular interest are scalar extensions of the SM that can change the nature of EWSB and the Higgs self-couplings. Indeed, there has been interest investigating scenarios in which the first evidence of new physics could be through variations of the Higgs tri-linear couplings DiLuzio:2017tfn ; Chang:2019vez ; Bahl:2022jnx ; Durieux:2022hbu . Another spectacular signature of additional scalars could be resonant di-Higgs production. Such a signal is much studied and appears in even the simplest of all possible extensions: a real scalar that is a singlet under the SM gauge group Silveira:1985rk ; OConnell:2006rsp ; Barger:2007im ; Bowen:2007ia ; Profumo:2007wc ; No:2013wsa ; Pruna:2013bma ; Profumo:2014opa ; Chen:2014ask ; Dawson:2015haa ; Buttazzo:2015bka ; Robens:2015gla ; Robens:2016xkb ; Kotwal:2016tex ; Lewis:2017dme ; Huang:2017jws ; Dawson:2018dcd ; Li:2019tfd ; Alves:2020bpi ; Papaefstathiou:2020iag ; Dawson:2020oco ; Adhikari:2020vqo ; Papaefstathiou:2021glr ; Dawson:2021jcl ; Hammad:2023sbd .
@@
There has been recent interest in moving beyond di-Higgs production and studying more generic di-scalar production. For example, non-resonant production of two new scalars can provide insight into the early universe electroweak phase transition Curtin:2014jma ; Chen:2017qcz . A more spectacular signal is asymmetric resonant production of the SM Higgs+a new scalar or symmetric resonant production of two new scalars Basler:2018dac ; Abouabid:2021yvw ; Adhikari:2022yaa ; Dawson:2017jja ; Costa:2015llh ; Robens:2019kga ; Papaefstathiou:2020lyp . In this paper, we develop benchmark points for such signals in the SM extended by a complex scalar that is a singlet under the SM gauge group Coimbra:2013qq ; Costa:2014qga ; Ferreira:2016tcu ; Muhlleitner:2017dkd ; Basler:2019nas ; Muhlleitner:2020wwk ; Egle:2022wmq ; Egle:2023pbm ; Barger:2008jx ; Alexander-Nunneley:2010tyr ; Dawson:2017jja ; Adhikari:2022yaa ; Costa:2015llh ; Robens:2019kga ; Papaefstathiou:2020lyp ; Dawson:2023oce . Beyond providing interesting collider signatures of new scalars, the complex singlet extension can also help provide a strong first order electroweak phase transition necessary to generate the baryon asymmetry of the universe via electroweak baryogenesis Barger:2008jx ; Chao:2014ina ; Jiang:2015cwa ; Chiang:2017nmu ; Cheng:2018ajh ; Grzadkowski:2018nbc ; Chen:2019ebq ; Chiang:2019oms ; Cho:2021itv ; Cho:2022our ; Cho:2022zfg ; Zhang:2023mnu . This extension has been much studied in the scenarios where there are additional $\mathbb{Z}_{2}$ or $U(1)$ symmetries McDonald:1993ex ; Cerdeno:2006ha ; Barger:2008jx ; Coimbra:2013qq ; Costa:2014qga ; Costa:2015llh ; Ferreira:2016tcu ; Muhlleitner:2017dkd ; Basler:2019nas ; Muhlleitner:2020wwk ; Robens:2019kga ; Papaefstathiou:2020lyp ; Egle:2022wmq ; Egle:2023pbm . We will focus on the scenario in which there is no additional symmetry beyond the SM gauge group Dawson:2017jja ; Adhikari:2022yaa .
@@
If the mixing between $h_{3}$ and the SM Higgs boson is minuscule, the coupling between $h_{3}$ and SM fermions and gauge bosons will be negligible. Hence, the production rate of $h_{3}$ through gluon fusion or vector boson fusion, $\sigma(pp\rightarrow h_{3})=\sin^{2}\theta_{1}\sin^{2}\theta_{2}\sigma_{\rm SM}(pp\rightarrow h_{3})$, is very small. Indeed, for mixing angles $\sin\theta_{1}\leq 0.2$ and $\sin\theta_{2}\leq 0.01$, which, as we will see later, are reasonable ranges, then $\sigma(pp\rightarrow h_{3})\leq 4\times 10^{-6}\cdot\sigma_{\rm SM}(pp\rightarrow h_{3})$. This is more than an order of magnitude below the expected 95% CL direct search limits of the FCC-hhEuropeanStrategyforParticlePhysicsPreparatoryGroup:2019qin . In that case, the major production mode of $h_{3}$ would then be directly through these decays. Hence, it is possible that searches for resonant production of $h_{2}\rightarrow h_{1}h_{3}$ and $h_{2}\rightarrow h_{3}h_{3}$ could provide a discovery of two new real scalars at once, and could be the only feasible mechanisms to discover $h_{3}$ Dawson:2017jja .
@@ II Model
$V(\Phi,S_{c})$ $=$ $-\frac{\mu^{2}}{2}\Phi^{\dagger}\Phi+\frac{\lambda}{4}(\Phi^{\dagger}\Phi)^{2}+\frac{b_{2}}{2}|S_{c}|^{2}+\frac{d_{2}}{4}|S_{c}|^{4}+\frac{\delta_{2}}{2}\Phi^{\dagger}\Phi|S_{c}|^{2}$ (2)
@@
where the mass eigenstates are $h_{1},\,h_{2},\,$ and $h_{3}$ with masses $m_{1},\,m_{2},\,$ and $m_{3}$, respectively. We will associate $h_{1}$ with the observed scalar and set $m_{1}=125$ GeV ATLAS:2022vkf ; CMS:2022dwd . There is, in principle, a third $SO(3)$ rotation angle in Eq. (4), but it can be removed via a phase rotation of $S_{c}\rightarrow e^{i\,\theta_{3}}S_{c}$ Dawson:2017jja . Such a phase rotation is at most a redefinition of parameters in the scalar potential [Eq. (2)]. Hence, the third rotation angle can be removed with no physical effect. The free parameters of the model are then
@@
By requiring that $(\langle\Phi^{\dagger}\Phi\rangle,\langle S_{c}\rangle)=(v_{\rm EW}^{2}/2,0)$ be an extremum of the potential and going to the scalar mass basis, the Lagrangian parameters can be related to these free parameters according to the following equations:
@@
Note that the coefficient of Eq. (15) differs by a factor of 2 from the other partial widths due to the two final state scalars not being identical. The general expressions for the trilinears in Eq. (13) are complicated, but using Eqs. (2,12), the $\theta_{2}\rightarrow 0$ limits of the relevant trilinear couplings are
$\lambda_{112}$ $=$ $\sin\theta_{1}\frac{m_{2}^{2}}{v_{\rm EW}}\left[\cos^{2}\theta_{1}\left(1+2\frac{m_{1}^{2}}{m_{2}^{2}}\right)+\frac{1}{\sqrt{2}}\cos\theta_{1}\sin\theta_{1}\left(\textrm{Re}\left(e_{1}\right)+\textrm{Re}\left(e_{2}\right)\right)\frac{v_{\rm EW}}{m_{2}^{2}}\right.$
$\left.-\left(1-\frac{3}{2}\sin^{2}\theta_{1}\right)\frac{v_{\rm EW}^{2}}{m_{2}^{2}}\left(\delta_{2}+\textrm{Re}\left(\delta_{3}\right)\right)\right]$
@@ III.1.1 Boundedness and Global Minima
$\frac{\lambda}{4}(\Phi^{\dagger}\Phi)^{2}+\frac{d_{2}}{2}|S_{c}|^{4}+\frac{\delta_{2}}{2}\Phi^{\dagger}\Phi|S_{c}|^{2}+\left(\frac{\delta_{3}}{4}\Phi^{\dagger}\Phi S_{c}^{2}+\frac{d_{1}}{8}S^{4}_{c}+\frac{d_{3}}{8}S_{c}^{2}|S_{c}|^{2}+{\rm h.c.}\right)\geq 0.$ (22)
@@ III.1.2 Perturbative Unitarity
Notably, the bound on $\lambda$ in Eq. (26) places a constraint on the allowed masses and mixings. For a particular choice of mixing angles, this induces an upper limit on the allowed heavy Higgs masses. In practice, as with the boundedness and global minimum constraints, we check numerically that the eigenvalues of Eq. (25) are bounded by $8\pi$.
@@ III.2 Experimental Constraints
We now consider experimental constraints on this model. First, we take Higgs precision measurements into account ATLAS:2022vkf ; CMS:2022dwd ; ATLAS-CONF-2022-067 ; CMS-PAS-HIG-19-011 . These measurements are typically given in terms of signal strengths, i.e. ratios of beyond SM (BSM) and SM predictions for different Higgs production and decay channels:
@@
Note that these conclusions are independent of the choice of the other mixing angle $\theta_{2}$. The mixing angle $\theta_{1}$ can then be fit using a $\chi^{2}$ distribution:
@@
where $\Gamma_{\rm SM}(h_{2})$ is the total SM-like width at the mass $m_{2}$. From this it can be shown that the branching ratio of $h_{2}$ into SM final states is
@@
Traditionally, a parameter point is accepted if the predicted cross section is lower than all of the observed 95% upper limits on cross sections reported in searches for new particles. The point is rejected if the predicted cross section exceeds an observed cross section limit in any channel. For example, see Refs. Bechtle:2013xfa ; Bechtle:2020uwn . We refer to this as the “hard cut” method. The red dashed curves in Figs. 1(a,b,c) show the resulting limits on $\sin\theta_{1}$ from utilizing the hard cut method. The list of included experimental results are given in Table 1 in App. D. Since $h_{2}$ couples to SM fermions and gauge bosons like the SM Higgs except suppressed by $\sin\theta_{1}$, its production and decay are similar to a SM Higgs of the same mass, except when the additional decay to scalars is available. Therefore, its dominate decays will be to gauge bosons or scalars and we only consider those experimental limits.
Figure 1: The 95% CL upper limits on $|\sin\theta_{1}|$ for (a) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$, (b) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, (c) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ profiled over, and (d) a variety of ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$. In all plots (black solid) the bounds from Higgs precision are shown. In (a-c) the bounds resulting from the (red dashed) hard cuts on scalar searches, (magenta dot-dot-dashed) $\Delta\chi^{2}$ scalar search fits from Eq. (40), and (blue solid) $\Delta\chi^{2}$ combination of scalar searches and Higgs precision from Eq. (41). The results in (d) are from the $\Delta\chi^{2}$ combination of scalar searches and Higgs precision.
@@
where $\sigma_{i}(pp\rightarrow h_{2})$ is the resonance production cross section from initial state $i$, ${\rm BR}(h_{2}\rightarrow f)$ is the branching ratio into final state $f$, and $\hat{\sigma}_{i,Exp}^{f}$ ($\hat{\sigma}_{i,Obs}^{f}$) is the expected (observed) 95%CL upper limit on $\sigma(i\rightarrow h_{2}){\rm BR}(h_{2}\rightarrow f)$. It can be checked that, for an individual search, this $\chi^{2}$ reproduces the limit $\sigma(i\rightarrow h_{2}){\rm BR}(h_{2}\rightarrow f)<\hat{\sigma}_{i,Obs}^{f}$, consistent with the hard cut method. With this definition, all heavy scalar search channels can be consistently combined:
@@
$\chi^{2}_{\rm Tot}=\chi^{2}_{h_{1}}+\chi^{2}_{h_{2}}.$ (41)
The magenta dot-dot-dashed lines in Figs. 1(a,b,c) show the 95% CL upper limits on $|\sin\theta_{1}|$ that result from fitting scalar searches according to Eq. (40). For all limits we have assumed ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})={\rm BR}(h_{2}\rightarrow h_{3}h_{3})=0$. In (a) we set ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$, in (b) ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, and in (c) we profile over ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$. As can be clearly seen, the results of Eq. (40) have similar behavior to and are consistently stronger than the hard cuts. This can be understood by noting that since all production cross sections of $h_{2}$ are proportional to $\sin^{2}\theta_{1}$, any increase in $|\sin\theta_{1}|$ essentially increases all signal rates in all channels uniformly increasing the $\chi^{2}$ value. Accordingly, we would more-or-less expect to see a global upward fluctuation in the channels listed in Tab. 1 if the scalar resonance exists. Such a global increase should result in a stronger constraint. Our $\chi^{2}$ method accounts for this, while the hard cut method does not.
The result from combining the scalar searches and Higgs precision data from Eq. (41) are shown as blue solid lines in Figs. 1 (a,b,c). For ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$, the combined $\chi^{2}_{\rm Tot}$ fit gives very similar results as the scalar search $\chi^{2}_{h_{2}}$. For the results with ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$ or ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ profiled over, combining the Higgs precision constraints with the scalar searches strengthens the bounds in the large mass region: $m_{2}\gtrsim 800$ GeV. Indeed, although the scalar search constraints may get weaker than the Higgs precision at very high mass, the combined Higgs and scalar search results stay more or similarly stringent as the Higgs measurement results.
@@
There are also constraints on the singlet mixing angle from electroweak precision observables Lopez-Val:2014jva ; deBlas:2016ojx ; Dawson:2017jja ; Ilnicka:2018def . For small $\theta_{2}$, fits to the oblique parameters Dawson:2017jja ; Ilnicka:2018def are always weaker than the most stringent constraints in Fig. 1. The strongest electroweak precision constraint comes from fitting the $W$-mass individually Lopez-Val:2014jva . This is most relevant in the high mass region where the $W$-mass limit is $|\sin\theta_{1}|\lesssim 0.20-0.21$ for $m_{2}\gtrsim 800$ GeV Lopez-Val:2014jva ; Dawson:2017jja . Again, the most stringent constraints in Fig. 1 are as strong or stronger than the $W$-mass limit. Hence, for the rest of the paper we consider constraints on $\sin\theta_{1}$ from direct searches and Higgs precision data only.
Figure 2: Maximum allowed resonant production rates for $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in (a) gluon fusion and (b) VBF channels for (red dashed) hard cut bounds, (magenta dot-dot-dashed) $\Delta\chi^{2}$ bounds on scalar searches from Eq. (40), and (blue solid) $\Delta\chi^{2}$ bounds on scalar searches and Higgs precision from Eq. (41). For comparison, SM predictions for (black solid) $h_{2}$ production and (green dot-dash-dashed) $h_{1}h_{1}$ production are shown.
In Fig. 2 we show 95% CL upper limits on the cross section for production and decay $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in both the (a) gluon fusion and (b) VBF channels. These bounds are found by performing a simultaneous fit to $\sin\theta_{1}$ and $\sin^{2}\theta_{1}{\rm BR}(h_{2}\rightarrow h_{1}h_{1}$) with ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})={\rm BR}(h_{2}\rightarrow h_{3}h_{3})=0$. For the $\Delta\chi^{2}$ fits, $\sin\theta_{1}$ is profiled over. For the hard cut, a value of $\sin^{2}\theta_{1}{\rm BR}(h_{2}\rightarrow h_{1}h_{1})$ is accepted if there exists a $\sin\theta_{1}$ that passes all cuts and Higgs precision data. For comparison, the SM single $h_{2}$ and double $h_{1}h_{1}$ production rates are shown. The $\Delta\chi^{2}$ and hard cut bounds produce very similar limits on $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$. For $m_{2}\lesssim 500$ GeV, the resonant gluon fusion di-Higgs rate can still be above the SM di-Higgs rate rate, up to an order of magnitude larger. For VBF, the allowed rate is above the SM VBF di-Higgs rate for all masses shown. The resonant VBF rate can be up to a factor of 20 larger than the SM VBF di-Higgs prediction.